# MaskBit: Embedding-free Image Generation via Bit Tokens

### v1

@@ B.1 Stage-I
For data augmentation, we closely follow the prior arts [55, 56]. Specifically, we use random cropping of 80%-100% of the original image with a random aspect ratio of 0.75-1.33. After that, we resize the input to $256\times 256$ and apply horizontal flipping.
@@
  • Perceptual loss weight: 0.7
@@
  • Optimizer: AdamW [37]
@@ B.2 Stage-II
For data augmentation, we closely follow the prior arts [55, 56]. Specifically, we use random cropping of 80%-100% of the original image. After that, we resize the input to $256\times 256$ and apply horizontal flipping.
@@
  • Optimizer: AdamW [37]
@@ B.3 Sampling Procedure
Our sampling procedure closely follows prior works for non-autoregressive image generation [13, 4, 56, 55, 5]. By default, the Stage-II network generates images in 64 steps. Each step, a number of tokens is unmasked and kept. The number of tokens to keep is determined by an $\arccos$ scheduler [3, 4]. Additionally, we use classifier-free guidance and a guidance scheduler [13, 55]. Specifically, for models trained on 256 resolution, we use a temperature of 7.3, a guidance of 7.6 and a “scale_pow”of 3.5.
@@ B.4 Sensitivity to Sampling Parameters
Appendix C More Discussion of Bit Tokens
@@
Appendix D More Reconstruction Results of Flipping Bit Tokens
@@
Appendix E Limitations
In this study, the focus has been on non-autoregressive, token-based transformer models for generation. Due to the computational costs, we did not explore using the modernized VQGAN model (VQGAN+) in conjunction with diffusion models. For similar reasons, MaskBit was only trained on ImageNet, thus inheriting the dataset bias, i.e., focus on a single foreground object. This makes it difficult to do have a qualitative comparison to methods trained on large-scale datasets such as LAION [45]. We also note, that the standard benchmark metric FID has received criticism due to it measuring only distribution similarities. However, due to the lack of better and established alternatives, this study uses FID for its quantitative evaluation and comparison.
Appendix F Broader Impacts

### v2

@@ B.1 Stage-I
For data augmentation, we closely follow the prior arts (Yu et al., 2023; 2024a). Specifically, we use random cropping of 80%-100% of the original image with a random aspect ratio of 0.75-1.33. After that, we resize the input to $256\times 256$ and apply horizontal flipping.
@@
  • Perceptual loss weight: 0.1
@@
  • Optimizer: AdamW (Loshchilov & Hutter, 2019)
@@ B.2 Stage-II
For data augmentation, we closely follow the prior arts (Yu et al., 2023; 2024a). Specifically, we use random cropping of 80%-100% of the original image. After that, we resize the input to $256\times 256$ and apply horizontal flipping.
@@
  • Optimizer: AdamW (Loshchilov & Hutter, 2019)
@@ B.3 Sampling Procedure
Our sampling procedure closely follows prior works for non-autoregressive image generation (Gao et al., 2023; Chang et al., 2022; Yu et al., 2024a; 2023; Chang et al., 2023). By default, the Stage-II network generates images in 64 steps. Each step, a number of tokens is unmasked and kept. The number of tokens to keep is determined by an $\arccos$ scheduler (Besnier & Chen, 2023; Chang et al., 2022). Additionally, we use classifier-free guidance and a guidance scheduler (Gao et al., 2023; Yu et al., 2023). We refer to the codebase available on https://github.com/markweberdev/maskbit/ for the detailed sampling parameters for each model.
@@ B.4 Sensitivity to Sampling Parameters
Appendix C Visualization of Masking Groups of Bit Tokens
Fig. 6 provides a visualization of the approach, masking groups of bit tokens, presented in this paper. When using a single group, a bit token will either be masked completely or remain unchanged after the masking is applied. When using more than a single group, the bit token is split into groups of consecutive bits. Each of these groups can be independently masked during the masking procedure, leading to some cases where the bit tokens are partially masked as shown in the illustration. Notably, within each group of bits, the masking can only be applied to all bits in the group.
Figure 6: Visualization of the process for masking of groups of bit tokens. The figure visualizes the masking procedure of bit tokens (12 bits) when using either one group (top row) or two groups (bottom row). When using a single group, a bit token can either be fully masked or fully non-masked. When using two groups, each bit token is split into two consecutive groups (i.e., the first group contains the first 6 bits and the second group has the last 6 bits). These two groups can be independently masked. This leads to the case where a bit token can be partially masked, as shown in the illustration.
Appendix D More Discussion of Bit Tokens
D.1 Connection of Semantic Structure and Generator Performance
Prior work (Gu et al., 2024) studying the tokenizer objective has found that a strong semantic representation in the latent space is important for good generation performance. This observation was made by using different variations of the LPIPS perceptual loss. The experiments presented in this work in Sec. 2 support the same observation. Specifically, VQGAN+ uses a ResNet logits-based perceptual loss which we found to work well for generation. However, as mentioned in the paragraph “Perceptual Loss”, adding intermediate activations of ResNet to the perceptual loss decreased the generation performance even though it helps the reconstruction performance.
Given the insights on the importance of semantics, this work studies the latent space of bit tokens and the connection to the image content. The key finding is that bit tokens exhibit this uniquely constrained structure by design, which leads to a structured semantic representation. In the next paragraph, we discuss how this structured representation affects the internal generation model feature representation per token. We refer to the experimental results presented in Sec. 4.1 for a quantitative study on how much benefit bit tokens with this structured semantic representation bring to the generation performance.
D.2 Implications of Different Token Spaces on the Generator
@@
Appendix E More Reconstruction Results of Flipping Bit Tokens
@@
Furthermore, we repeat this experiment in a zero-shot manner on COCO (Lin et al., 2014) with the same model only trained on ImageNet. The results are visualized in Fig. 8, demonstrating that the bit token’s structured semantic representation also holds for COCO images.
Appendix F Zero-Shot Reconstruction Results on COCO
In Tab. 5, we provide zero-shot evaluation of the presented tokenizers of Sec. 2 on COCO (Lin et al., 2014) to show how the modifications made to the orginal VQGAN (Esser et al., 2021) impact the performance on more complex datasets. For this experiment, we use the same models as for the experiments in Sec. 2 that were only trained on ImageNet (Deng et al., 2009). All images are resized with their shorter edge to 256 and then center cropped before feeding them into the models. The results show that the modifications made to obtain VQGAN+ and its embedding-free variant also achieve improvements on more complex datasets in the zero-shot evaluation setting.
Table 5: Zero-shot reconstruction evaluation on COCO with $256\times 256$ resolution.
Model zero-shot rFID $\downarrow$
Taming-VQGAN (Esser et al., 2021) 16.3
VQGAN+ 8.7
Embedding-free variant 8.3
Appendix G Nearest Neighbor Analysis of Generated Images
Following the idea presented in Esser et al. (2021), we conduct a nearest neighbor analysis. In this analysis, a generated image is used and compared to all images in the training set. For this comparison, we use two distance measures: LPIPS and Hamming distance. The LPIPS metric is used to measure perceptual similarity of two images, while the Hamming distance takes the bit tokens of the encoded images and computes the bit-wise distance. In Fig. 9 and Fig. 10, we visualize the 10 nearest neighbors for each generated image and distance measure. The results show that the model is not just memorizing samples from the training set, but generates indeed new images. Since LPIPS and Hamming distance are two different measures, the obtained nearest neighbors also differ. For example, considering the neighbors to the generated images of the parrot or the space rocket in Fig. 9, we observe that the nearest samples according to LPIPS focus more on staying true to the reference color palette, while this is less the case for the closest samples according to the Hamming distance. This finding aligns with the observations in the bit-flipping experiment, where with small changes in the Hamming distance, the visual attributes can vary quite a bit, while the semantic content stays the same.
Appendix H Limitations
In this study, the focus has been on non-autoregressive, token-based transformer models for generation. Due to the computational costs, we did not explore using the modernized VQGAN model (VQGAN+) in conjunction with diffusion models. For similar reasons, MaskBit was only trained on ImageNet, thus inheriting the dataset bias, i.e., focus on a single foreground object. This makes it difficult to do have a qualitative comparison to methods trained on large-scale datasets such as LAION (Schuhmann et al., 2022). We also note, that the standard benchmark metric FID has received criticism due to it measuring only distribution similarities. However, due to the lack of better and established alternatives, this study uses FID for its quantitative evaluation and comparison.
Appendix I Broader Impacts
@@
Figure 8: Examples demonstrating the structured representation of bit tokens in a zero-shot setting. The structured semantic representation can also be found when running the tokenizer in a zero-shot manner on the more complex dataset COCO (Lin et al., 2014).
Figure 9: Nearest Neighbors of two generated Images. The figure shows the ten nearest neighbors in the ImageNet training set in terms of LPIPS (top) and Hamming distance (bottom).
Figure 10: Nearest Neighbors of two generated Images. The figure shows the ten nearest neighbors in the ImageNet training set in terms of LPIPS (top) and Hamming distance (bottom).